b)  Answer the following stoichiometry questions by referring to the equation below:

2

KClO3               2 KCl + 3 O2

i.        If 1.50 moles of KClO3 decomposes, what is the mass of O2 that will be produced?

ii.       If 80.0 grams of O2 was produced, how many moles of KClO3 are decomposed?

iii.    Find the mass of KClO3 needed if we need to produce 2.75 moles of KCl.

1
Expert's answer
2022-05-16T17:05:04-0400

I. moles of O2 produced are 4.5 moles, M(O2) = 32, m(O2) = 32 × 4.5= 144 g.

II. n = m/M = 122.6/80 = 1,53 mol

III. moles of 2.75/2 = 1.375moles, m = 1.375× 122.6 = 168.6 g.